Part B A cylinder with a 16-cm-diameter piston contains gas at a pressure of 3.0 atm How much force does the environment exert on the piston? 2040 SubmitP Previous Answers Request Answer X Incorrect; Try Again; 9 attempts remaining Part The gas expands at constant pressure and pushes the piston out 10 cm. How much work is done by the environment? Submit Request Answer
Answer #1

here,

diameter , d = 16 cm

radius , r = d/2 = 8 cm = 0.08 m

pressure , P = 3 atm = 3 * 1.01 * 10^5 Pa

B)

the force required , F = P * pi * r^2

F = 3.03 * 10^5 * pi * 0.08^2 N

F = 6.09 * 10^3 N

C)

x = 10 cm = 0.1 m

the work done , W = F * x

W = 6.09 * 10^2 J